b. Average variable cost, average fixed cost, marginal cost and average cost are necessary for the analysis of the cost structure of the firm. Define each of these terms and show how they are calculated

c. Describe the relationship between the marginal product and the total product of a firm
Answer #1

Average variable cost can be obtained by dividing total variable cost by number of units of output. Average variable cost isAFC = average fixed cost TFC = total fixed cost Q = number of units Example: TFC = 1000 Q = 10 AFC = 1000/10 AFC = 100 AveragExample: TC = 2500 Q = 10 ATC = 2500/10 ATC = 250 Or ATC = AFC + AVC ATC = 100 + 150 ATC = 250 Marginal cost is the additionan=number of units of output Example: If Q = 10, TC = 2500 If Q = 11, TC = 2700 Thus, MC will be MC = 2700 - 2500 MC = 200 TotGraph: Output TP maximum MP-0 Labor units MP